a person climbs up a stalled escalator in 60s. if standing on the same but escalator running with constant velocity he takes 40s. how much time is taken by the person to walk up the moving escalator/

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Assume the length of the escalator is 60 ft
then
Walking: 1 ft/sec
moving: 60/40 = 1.5 ft/sec
walking & moving 2.5 ft/sec
:
Find the time at 2.5 ft/sec
t = 60/2.5 = 24 seconds walking & moving
 
Regards
Arun(askIITians forum expert)